If a buffer solution is 0.510 M in a weak acid (K, = 4.1 x 10-6) and 0.300 M in its conjugate base, what is the pH? pH =
Answer #1


pH = pka+log (Salt) (acid] 0.510 M, Conjugate love = salt = 0.300M . pka = -log (ka) -log(klx106) 5.39 PH= 539+10g (0:30) = 5